OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 6) and a directrix at y = -6.

OpenStudy (anonymous):

here the vertex is origin hence the required parabola is x^2= 4*6*y x^2=24y

OpenStudy (anonymous):

well, another way of doing it is: \[\sqrt{x ^{2}+(y-6)^{2}} =0+y+6/1\], so, x^2=24y I have equated the perpendicular distance from a point P(x,y) on the directrix y+6=0, and the distance of point P from the focus.....from the basic definition of parabola, these two distances should always be equal...

OpenStudy (anonymous):

thank you i recognize the second way more. :)

OpenStudy (anonymous):

U are welcome!